# PersonaCraft: Personalized Full-Body Image Synthesis for Multiple Identities from Single References Using 3D-Model-Conditioned Diffusion

### v1

@@ Abstract
Personalized image generation has been significantly advanced, enabling the creation of highly realistic and customized images. However, existing methods often struggle with generating images of multiple people due to occlusions and fail to accurately personalize full-body shapes. In this paper, we propose PersonaCraft, a novel approach that combines diffusion models with 3D human modeling to address these limitations. Our method effectively manages occlusions by incorporating 3D-aware pose conditioning with SMPLx-ControlNet and accurately personalizes human full-body shapes through SMPLx fitting. Additionally, PersonaCraft enables user-defined body shape adjustments, adding flexibility for individual body customization. Experimental results demonstrate the superior performance of PersonaCraft in generating high-quality, realistic images of multiple individuals while resolving occlusion issues, thus establishing a new standard for multi-person personalized image synthesis. Project page: https://gwang-kim.github.io/persona_craft
Figure 1: PersonaCraft generates realistic, personalized images of multiple people with complex occlusions, preserving full body details, including facial identity and body shape from single references. It also supports a novel application: user-defined body shape control, allowing custom body shape adjustments.
@@ 1 Introduction
Recent advances in personalized image generation [20, 31, 33, 60, 61, 65, 17, 4, 23, 69, 70, 56, 82, 83, 10, 18, 25, 52, 64, 85, 74, 24, 71, 48, 78, 19, 84, 64, 80] have enabled the creation of visually compelling and highly customized portraits. Diffusion models [34, 66, 84, 62, 14, 42, 39, 41, 40] have been especially impactful, significantly enhancing the adaptability and quality of generative methods through improved controllability, such as pose-guided generation with ControlNet [81]. However, generating images with multiple individuals remains challenging in spite of its development [11, 30, 50, 26, 46, 28, 58, 75, 43], particularly in complex scenarios with occlusions, where overlapping body parts or interacting poses obscure person-specific details (yellow boxes in the baselines in Fig. 1(a)).
Full-body personalization also poses significant challenges, as most existing methods primarily focus on facial identity, often overlooking individual body shape details. Many approaches rely on 2D pose estimations [79, 16, 29], which lack the depth required to capture unique physical attributes accurately. Furthermore, current methods struggle to effectively separate a person’s identity from their clothing, resulting in unwanted style transfers and inconsistencies, particularly during occluded interactions or when depicting intricate garments.
To address these limitations, we propose PersonaCraft, a framework for generating highly personalized, full-body images of multiple individuals in complex scenes. PersonaCraft integrates diffusion models with 3D human modeling (SMPLx model [57]) for precise control over body shape and pose. The approach includes three core mechanisms: 1) Face and body identity extraction: We extract body shape parameters via SMPLx fitting [13, 21, 2] and obtain face ID embeddings from reference images using InsightFace [1]. 2) 3D-aware pose conditioning with SMPLx-ControlNet (SCNet): SCNet, which leverages SMPLx depth maps and is trained on our curated MPII-SMPLx dataset featuring complex, occluded multi-human scenes, enables robust control over body shape and pose, allowing realistic interactions between individuals. 3) Multiple Human Personalized Image Synthesis: For the personalization of multiple identities, PersonaCraft uses seamlessly SCNet and IdentityNet [71], leveraging face masks to accurately localize identities. Unlike two-stage methods (e.g., OMG [43]) or test-time optimization approaches (e.g., DreamBooth [60]), PersonaCraft efficiently personalizes multiple identities within a single generative process.
PersonaCraft generates highly realistic and personalized images of people, even in occluded scenarios. As shown in Fig. 1(a), prior methods struggle with body shape personalization and anatomical inaccuracies. In contrast, PersonaCraft ensures accurate body shapes and face identities while maintaining distinct styles.
Moreover, PersonaCraft introduces a unique feature: User-defined body shape control, allowing users to adjust attributes like height and body build for tailored body shapes as shown in Fig. 1(b). This feature is valuable for applications in advertising and social media, where customization boosts engagement.
Extensive quantitative and qualitative evaluations, including user studies, validate PersonaCraft’s superior performance in preserving facial identity, body shape, and natural human anatomy, even in challenging scenarios involving multiple people and significant occlusions. These capabilities position PersonaCraft as a state-of-the-art solution for high-quality, multi-person image generation, surpassing existing methods in accuracy and personalization.
Figure 2: Overview of PersonaCraft. We extract face and body embeddings from reference images, then perform personalized image synthesis using controllable modules for face identity and body shape customization. Our method applies 3D-aware pose conditioning with SMPLx-ControlNet (SCNet) to generate occlusion-robust images that accurately reflect body shapes.
@@ 2 Related Works
Single-concept personalization. Text-to-image (T2I) diffusion models and applications [34, 66, 84, 62, 14, 42, 39, 41, 40] enable single-concept personalization, adapting pre-trained models for individual subjects. Early methods used optimization-based techniques [20, 31, 33, 60, 61, 65, 17] or textual embeddings [4, 23, 69, 70, 56, 82, 83]. LoRA methods [35, 68] reduced the need for many trainable parameters. Recent works[10, 18, 25, 52, 64, 85, 74, 24, 71, 48, 78, 19, 84, 64, 80] like IP-Adapter [80] and InstantID [71] adopts modular design enabling fast personalization from a single reference image and included human-centric modules. However, these methods often struggle with full-body fidelity, either neglecting body shape or failing to disentangle identity from clothing in the case of single reference personalization.
Multi-concept personalization. Current approaches often combine multi-concept datasets through joint training, adding losses or optimization steps to facilitate model merging. Some methods [11, 30, 50, 26] use cross-attention maps to avoid entangling concepts. Custom Diffusion [46] proposes either joint training or constrained optimization for model integration. Mix-of-show [28] utilize gradient fusion reduces identity loss, while regionally controllable sampling manages attribute binding. Modular Customization [58] leverages orthogonal directions to isolate concepts for easier integration. FastComposer [75] enhances text conditioning in diffusion models with subject embeddings, allowing personalization without retraining. OMG [43] proposes the method without additional training for multi-concept personalization enhancing efficiency and flexibility. However, existing methods still exhibit limitations in maintaining occlusion robustness and precise control over body shapes.
Pose-guidance for image synthesis. Recent works such as ControlNet [81] and HyperHuman [49] combine T2I diffusion with 2D skeleton poses for controllable human image synthesis. DWpose [79] improves on OpenPose [16] by providing more accurate skeletons to enhance diffusion models’ quality [22, 36], while DensePose [29] maps correspondences between RGB images and 3D surfaces. PODIA-3D [41] uses depth maps for pose-guidnace but not for the human image synthesis. The SMPL [51] and SMPLx [57] model , widely used for realistic human body and shape modeling, provide detailed 3D pose and body shape parameters. Champ [86] employs SMPL as a 3D condition for video synthesis but is limited to single-human cases.
Unlike previous methods, our approach uniquely handles full-body personalization with occlusion robustness and 3D body shape conditioning. Using SMPLx ControlNet (SCNet), we synthesize realistic full-body images that adapt to complex poses and occlusions, preserving face and body identity. Our method integrates seamlessly with single-reference images, enabling efficient personalization. Additionally, it offers user control over body shapes, creating a versatile human-centric generation framework that addresses limitations in existing T2I personalization techniques.
@@ 3 Proposed Method: PersonaCraft
Figure 3: SMPLx-ControlNet (SCNet) leverages 3D-aware pose conditioning using SMPLx models to accurately capture body shape and pose. The SMPLx depth map, generated from body shape parameters and 3D poses, serves as a precise conditioning signal for robust image synthesis, effectively handling occlusions.
We aim to generate high-quality, personalized images of multiple people in a complex scene from the single full-body references. Given a set of full-body reference images $\{I_{\text{ref}}^{(i)}\}^{N_{\text{human}}}_{i=1}$, target poses $\{\bm{p}^{(i)}\}^{N_{\text{human}}}_{i=1}$, where $N_{\text{human}}$ is the number of human identities, and a text prompt $y$, our goal is to synthesize a personalized image $I_{P}$ that accurately reflects the identities, poses, and the semantic content specified by the text prompt. This task is particularly challenging due to complex occlusions and the need for precise full-body shape personalization.
Our proposed method, PersonaCraft, tackles these challenges using a 3D human model-based diffusion model, as shown in Fig. 2. We begin by extracting face and body identity embeddings from the reference images. Next, we apply a 3D-aware pose conditioning technique with SMPLx-ControlNet (SCNet) to generate high-quality, occlusion-robust images. We then perform personalized image synthesis for multiple humans. Finally, we introduce a novel application of our method: user-defined body shape control, allowing users to specify desired body shapes and generate customized human images.
3.1 Face and Body Identity Extraction
We first extract face and body identity information from a single reference image $I_{\text{ref}}^{(i)}$ per individual as illustrated in the left part of Fig. 2. We employ MultiHMR [13], a state-of-the-art SMPLx [57] fitting method, to perform robust 3D human reconstruction, estimating SMPLx parameters from a single image. This process extracts body shape parameters $\bm{\beta}^{(i)}$ for each person, ensuring accurate shape representation. For face identity, we utilize InstantID [71]’s face ID embedding method via InsightFace [1], obtaining face embeddings $\bm{f}^{(i)}$. These combined embeddings, $\bm{\beta}^{(i)}$ and $\bm{f}^{(i)}$, capture person-specific identity features, enabling the generation of highly personalized multi-human images.
Figure 4: Qualitative comparison of our method with the baselines demonstrates PersonaCraft’s superior performance in multi-identity images with occlusions. Our approach maintains consistent identities and accurate body shapes across varying poses, outperforming the baseline in handling occlusions, preserving body shape, and ensuring pose fidelity, even as the number of subjects increases.
3.2 3D-Aware Pose Conditioning using SMPLx-ControlNet
Existing human personalization methods often utilize 2D skeleton-based pose ControlNets [81], for pose-guided image generation. However, 2D pose estimation lacks the depth and specificity required to accurately capture unique body contours and structures, often leading to a loss of individual physical attributes. This limitation hinders full human personalization, including both face identity and body shape, and makes the models vulnerable to occlusions.
To address these issues, we leverage 3D pose information, $\bm{p}\in{N_{\text{joint}}\times 3}$, along with camera view pose $\bm{c}$, where $N_{\text{joint}}$ is the number of joints. 3D pose provides a more comprehensive representation of human pose, capturing depth information which can be significant in scenarios with complex poses and occlusions. However, even with 3D pose, naively using 3D skeleton-based pose conditioning may not fully capture body shape variations.
To overcome this limitation, we propose a novel 3D-aware pose conditioning technique using SMPLx-ControlNet (SCNet). By leveraging 3D human models, specifically SMPLx [57], we can accurately represent body shape and pose. Given body shape parameters $\{\bm{\beta}^{(i)}\}^{N_{\text{human}}}_{i=1}$, 3D poses $\{\bm{p}^{(i)}\}^{N_{\text{human}}}_{i=1}$, and a camera view pose $\bm{c}$, we first generate a 3D human model using SMPLx. We then render the SMPLx depth map $\bm{d}_{\text{SMPLx}}$, which serves as a strong conditioning signal for the diffusion model and effectively handles occlusions. This SMPLx depth map is used as a precise guidance for image generation with SCNet.
MPII-SMPLx Dataset. We train SCNet by fine-tuning depth ControlNet [81], originally trained on a large text-image-depth triplet dataset, using our curated MPII-SMPLx dataset. This dataset builds upon the MPII dataset [9], a comprehensive collection of human activity images featuring diverse and challenging poses and complex interactions. Notably, there is currently no well-curated multi-human image dataset with 3D human reconstruction, making our MPII-SMPLx dataset one of our key contribution. To enhance the MPII dataset, we incorporate textual descriptions from Khan et al.[38]. For 3D human reconstruction, we employ MultiHMR[13], a state-of-the-art multi-human SMPLx [57] fitting method, to handle multiple identities per image and render depth maps from the reconstructed SMPLx meshes. To account for occlusion scenarios, we balance the dataset with a 1:2 ratio of single-person to multi-person images. Depth clipping is applied during depth rendering to retain only values below 5, ensuring consistent quality. After preprocessing, we curate a final dataset of 6,348 image-text-SMPLx-parameter pairs. This carefully curated dataset enables robust model training with diverse 3D human poses, complex interactions, and detailed human parameters such as body shape and pose conditioning.
Training objective of SCNet. Following the ControlNet [81] framework, the training objective minimizes the difference between generated and ground truth images. Given an input image $\bm{x}_{0}$, its noisy counterpart $\bm{x}_{t}$ from noise $\epsilon$ at diffusion timestep $t$, a text prompt $y$, condition $\bm{c}$, and control signal $\bm{d}_{\text{SMPLx}}$, the SCNet $\mathcal{E}^{\text{SC}}_{\phi}$ is fine-tuned from depth ControlNet using the objective:

### v2

# PersonaCraft: Personalized and Controllable Full-Body Multi-Human Scene Generation Using Occlusion-Aware 3D-Conditioned Diffusion

@@ Abstract
We present PersonaCraft, a framework for controllable and occlusion-robust full-body personalized image synthesis of multiple individuals in complex scenes. Current methods struggle with occlusion-heavy scenarios and complete body personalization, as 2D pose conditioning lacks 3D geometry, often leading to ambiguous occlusions and anatomical distortions, and many approaches focus solely on facial identity. In contrast, our PersonaCraft integrates diffusion models with 3D human modeling, employing SMPLx-ControlNet, to utilize 3D geometry like depth and normal maps for robust 3D-aware pose conditioning and enhanced anatomical coherence. To handle fine-grained occlusions, we propose Occlusion Boundary Enhancer Network that exploits depth edge signals with occlusion-focused training, and Occlusion-Aware Classifier-Free Guidance strategy that selectively reinforces conditioning in occluded regions without affecting unoccluded areas. PersonaCraft can seamlessly be combined with Face Identity ControlNet, achieving full-body multi-human personalization and thus marking a significant advancement beyond prior approaches that concentrate only on facial identity. Our dual-pathway body shape representation with SMPLx-based shape parameters and textual refinement, enables precise full-body personalization and flexible user-defined body shape adjustments. Extensive quantitative experiments and user studies demonstrate that PersonaCraft significantly outperforms existing methods in generating high-quality, multi-person images with accurate personalization and robust occlusion handling.
Figure 1: PersonaCraft generates realistic, personalized images of multiple individuals with complex occlusions, preserving facial identity and body shape using occlusion-aware 3D pose and shape conditioned diffusion. PersonaCraft outperforms baselines in body shape personalization (blue arrows indicate failures) and naturalness (yellow arrows highlight artifacts, with zoomed-in views in yellow boxes).
@@ 1 Introduction
Recent advances in personalized image generation [21, 30, 32, 66, 67, 74, 18, 2, 22, 78, 79, 60, 92, 94, 8, 19, 24, 54, 73, 96, 84, 23, 81, 50, 88, 20, 95, 73, 90] and controllable image generation [91, 51, 28, 98, 37, 80, 16], powered by diffusion models [33, 75, 95, 69, 13, 42, 39, 41, 40], have paved the way for synthesizing multi-human scenes with explicitly defined poses. Despite these breakthroughs, generating images with multiple individuals in complex, occlusion-heavy environments remains a formidable challenge. Conventional approaches that rely on 2D skeleton-based conditioning lack the depth cues necessary to capture overlapping body parts and intricate inter-person interactions, often resulting in ambiguous occlusions and anatomical distortions. Although recent works [98, 16, 55] have begun exploring the use of 3D pose information, they are limited to single-person scenarios, thereby restricting their applicability to multi-person scenes.
Furthermore, achieving full-body personalization remains a significant challenge. Most existing methods prioritize facial identity, often overlooking the distinct characteristics of body shape. Although several approaches [90, 97] have enabled full-body personalization using general image encoders [65], they often struggle to disentangle intrinsic body shape from clothing attributes.
Figure 2: SMPLx-ControlNet (SCNet) with SMPLx depth and normal improves occlusion handling in pose generation compared to 2D OpenPose, though fine-grained occlusions remain challenging. Using our occlusion-focused methods, OccNet and OccCFG, we achieve superior pose consistency and anatomical coherence.
To address these challenges, we propose PersonaCraft, a framework for controllable and occlusion-robust full-body personalized image synthesis of multiple individuals in complex scenes. Firstly, our approach integrates diffusion models with 3D human modeling, employing SMPLx-ControlNet (SCNet), to utilize critical 3D geometric information to overcome the limitations of conventional 2D pose conditioning. Through careful exploration of 3D pose representations, we found that combining SMPLx depth and normal maps offers the most robust conditioning, ensuring high pose consistency and anatomical naturalness. However, despite the benefits of 3D-aware conditioning, handling complex, fine-grained occlusions still remains challenging, as shown in Fig. 2. Secondly, we propose the Occlusion Boundary Enhancer Network (OccNet), which explicitly enhances occluded regions at a finer level by leveraging depth edge signals that highlight occlusion boundaries. Complemented by our Occlusion-Aware Classifier-Free Guidance (OccCFG) strategy, our OccNet adaptively reinforces conditioning in occlusion-heavy areas without compromising contrast in unoccluded regions. Together, these components ensure robust anatomical coherence, even in complex multi-human interactions with significant occlusions, as presented in Fig. 2.
Another key advantage of our 3D-aware pose conditioning, beyond enhancing pose accuracy with improved occlusion handling and anatomical coherence, is its ability to encode body shape through SMPLx body shape coefficients. Our method can seamlessly be combined with Face Identity ControlNet [81] to achieve full-body multi-human personalization, representing a significant advancement beyond prior methods focused solely on facial identity [43, 97, 31, 93, 85, 81, 90]. Moreover, to overcome the limitations of SMPLx-based body representation, especially in challenging cases where body shape characteristics conflict, we introduce a dual-pathway approach where an optional textual pathway refines body shape. PersonaCraft enables fine-scale, user-defined body shape control through reference-based and interpolation/extrapolation-based adjustments.
Through extensive quantitative and qualitative evaluations, including user studies, our PersonaCraft shows significant improvements in occlusion handling, pose consistency, setting a new benchmark for anatomically precise controllable and personalized multi-human image synthesis.
Figure 3: Overview of our occlusion-aware 3D pose and shape conditioning. We generate SMPLx renderings for SMPLx-ControlNet (SCNet) and derive occlusion masks for the Occlusion Boundary Enhancer Network (OccNet) by counting intersected faces and masking depth edges. Combining SCNet and OccNet residuals with occlusion masks enables the base U-Net to handle occlusion-aware 3D pose and shape conditioning. Occlusion-aware classifier-free guidance (OccCFG) further improves anatomical coherence in occluded regions.
@@ 2 Related Works
Single-Concept Personalization. Text-to-image (T2I) diffusion models and applications [33, 75, 95, 69, 13, 42, 39, 41, 40] enable single-concept personalization, adapting pre-trained models for individual subjects. Early methods used optimization-based techniques [21, 30, 32, 66, 67, 74, 18] or textual embeddings [2, 22, 78, 79, 60, 92, 94]. LoRA methods [34, 77] reduced the need for many trainable parameters. Recent works [8, 19, 24, 54, 73, 96, 84, 23, 81, 50, 88, 20, 95, 73, 90] like IP-Adapter [90] and InstantID [81] adopt modular designs for fast personalization from a single reference image and human-centric modules. However, these methods face challenges in handling complex human poses, neglect body shape preservation, or fail to disentangle identity from clothing in personalized human image synthesis.
Multi-Concept Personalization. Recent multi-concept personalization methods [9, 29, 52, 25] use cross-attention to prevent concept entanglement. Custom Diffusion [46] integrates models via joint training or constrained optimization. Mix-of-Show [27] employs gradient fusion for identity preservation and regional sampling for attribute binding. Modular Customization [63] isolates concepts via orthogonal directions. FastComposer [85] enhances training-free personalization. OMG [43] propose 2-stage multi-concept personalization without training. StoryMaker [97] ensures multi-character consistency via segmentation-constrained cross-attention. UniPortrait [31] propose ID embeddings and routing modules for high fidelity and editability. ID-Patch [93] mitigates ID leakage and ensures precise identity-position association introducing ID patches. However, existing methods still exhibit limitations in maintaining occlusion robustness and precise control over body shapes.
Controllable Human Generation. Recent works, such as ControlNet [91] and HyperHuman [51], integrate T2I diffusion with 2D skeletons for controllable human synthesis. HumanSD [37] enhances skeleton-based generation with heatmap-guided denoising, while Stable-Pose [80] improves pose alignment using attention masking. DWpose [89] refines OpenPose [17] for better skeleton accuracy, and DensePose [28] maps RGB images to 3D surfaces. However, lacking depth cues, these methods struggle with occlusions in multi-human or complex pose scenarios. PODIA-3D [41] incorporates depth for pose guidance but not image synthesis. The SMPL [53] (Skinned Multi-Person Linear Model) and SMPLx [61] are parametric human body models that provide realistic 3D representations of human pose and shape. SMPL models the body using shape parameters and pose parameters with a differentiable blend skinning function, making it widely used in vision and graphics applications. SMPLx extends this by incorporating expressive facial and hand articulation, enabling more detailed human motion modeling. Recent works leverage SMPL for generative tasks: Champ [98] applies SMPL for single-human video synthesis, while HumanLDM [16] integrates SMPL with ControlNet for stable pose-conditioned human generation. However, these methods focus solely on single-human cases and do not address complex occlusions in multi-human scenarios.
@@ 3 Proposed Method: PersonaCraft
We propose PersonaCraft, a framework for controllable and occlusion-robust full-body personalized image synthesis of multiple individuals in complex scenes. PersonaCraft integrates diffusion models with 3D human modeling, leveraging SMPLx-ControlNet (SCNet) to incorporate 3D geometry, including depth and normal maps, for robust 3D-aware pose conditioning and enhanced anatomical coherence (Sec. 3.1). To handle fine-grained occlusions, we introduce the Occlusion Boundary Enhancer Network (OccNet), which exploits depth edge signals through occlusion-focused training, and the Occlusion-Aware Classifier-Free Guidance (OccCFG) strategy, which selectively reinforces conditioning in occluded regions without affecting unoccluded areas (Sec. 3.2). Furthermore, PersonaCraft seamlessly integrates with Face Identity ControlNet, enabling full-body multi-human personalization and surpassing prior approaches that focus solely on facial identity. Our dual-pathway body shape representation, combining SMPLx-based shape parameters with textual refinement, ensures precise full-body personalization and flexible user-defined body shape adjustments (Sec. 3.3).
Existing methods rely on 2D skeleton-based ControlNets [91], which lack depth cues and fail to capture identity-specific body shapes in multi-human scenes. We propose a 3D-aware pose conditioning technique, SMPLx-ControlNet (SCNet), that leverages SMPLx [61] to incorporate depth information, enabling occlusion-aware conditioning and body shape control.
Figure 4: Training of SMPLx-ControlNet (SCNet) and Occlusion Boundary Enhancer Network (OccNet). The networks are trained separately, with SMPLx depth, normal maps, and occlusion masks extracted from training images. The pretrained ControlNet [91] is fine-tuned with these 3D pose representations.
3.1 3D-Aware Pose and Shape Conditioning for Multi-Human Generation
Existing methods rely on 2D skeleton-based ControlNets [91], which lack depth cues and fail to capture identity-specific body shapes in multi-human scenes. We propose a 3D-aware pose conditioning technique using SMPLx-ControlNet (SCNet). By leveraging SMPLx [61], we represent pose with depth information, enabling occlusion and controlling body shape. As shown in Fig. 3, given 3D poses $\{\bm{p}^{(i)}\}_{i=1}^{N_{\text{human}}}$, body shape parameters $\{\bm{\beta}^{(i)}\}_{i=1}^{N_{\text{human}}}$, and camera parameters $\bm{c}$, we construct a 3D human scene of $N_{\text{human}}$ individuals with occlusions using SMPLx meshes. SMPLx renderings, like depth $\bm{d}_{\text{SMPLx}}$, serve as conditioning signals for the diffusion model, enabling high-fidelity image generation with improved identity preservation and occlusion handling.
As presented in Fig. 4, we train SCNet $\mathcal{E}^{\text{SC}}_{\phi}$ by leveraging the pretrained ControlNet [91] framework, with the following objective:
$\small\mathbb{E}_{\bm{x}_{0},t,y,\bm{d}_{\text{SMPLx}},\bm{\epsilon}}\Big{[}\Big{\|}\bm{\epsilon}-\bm{\epsilon}_{\theta}\Big{(}\bm{x}_{t},t,y,\mathcal{E}^{\text{SC}}_{\phi}(\bm{d}_{\text{SMPLx}})\Big{)}\Big{\|}_{2}^{2}\Big{]},$ (1)
where $\bm{\epsilon}\sim\mathcal{N}(0,1)$, $\bm{\epsilon}_{\theta}$ is a base text-to-image diffusion model, $t$ is the diffusion timesteps $y$ is the text prompt, $\bm{x}_{t}$ is the noisy latent representation of the image at timestep $t$, and it gets progressively refined during the denoising process. We evaluate different SMPLx renderings—depth, normal, and RGB—for conditioning and find that combining depth SCNet (SCNet-D, $\mathcal{E}^{\text{SC}}_{\phi_{D}}$, for occlusion cues) and normal SCNet (SCNet-N, $\mathcal{E}^{\text{SC}}_{\phi_{N}}$, for surface orientation) by integrating the residuals from each SCNet provides the most robust pose consistency and anatomical coherence (see Tab. 5). Therefore, we adopt depth-normal as the default 3D representation for SCNet.
Figure 5: Full pipeline of PersonaCraft for multi-human full-body personalization. By integrating our method with face embeddings from InsightFace [1] and Face Identity ControlNet [81], we advance toward full-body multi-human personalization.
3.2 Occlusion-Focused Generation
Although 3D-aware pose conditioning improves handling occluded regions, generating small, detailed occlusions remains difficult due to inter-person contact and self-occlusion, which cause anatomical distortions. To address this, we introduce occlusion masks from SMPLx-rendered 3D scenes, marking pixels as occluded if covered by multiple surfaces, as shown in Fig. 3 (right). For each pixel $(i,j)$, we count the number of faces $n_{\text{faces}}(i,j)$ that a projected ray passes through. A pixel is considered occluded if the ray intersects more than two surfaces. The occlusion mask $M_{\text{occ}}$ is defined as:
$\footnotesize M_{\text{occ}}(i,j)=\begin{cases}1,&\text{if }n_{\text{faces}}(i,j)>2\\ 0,&\text{otherwise}\end{cases}$ (2)
To reduce subtle occlusions and prevent overfitting, we filter small occlusions and refine the boundaries by dilating the mask’s edges. This improves the representation of occluded regions, especially in poses with inter-person contact and self-occlusion.
Figure 6: Qualitative comparison of personalized multi-human scene generation. Blue arrows indicate failures in body shape personalization, while yellow arrows highlight unnatural anatomical structures, with yellow boxes providing zoomed-in views.
Table 1: Quantitative evaluation of multi-human personalization across face identity, body shape preservation, pose accuracy, text alignment, and image quality. ‘Single’ denotes personalization for a single individual, while ‘Multi’ refers to cases with multiple identities (2–5), with ‘Total’ representing the averaged results.
Multi-Human Personalization Face ID preservation↑ Body shape preservation↑ Pose Text Image quality Single Multi Total Single Multi Total MPJPE (3D)↓ AP-0.5 (2D)↑ CLIP sim ↑ IS ↑ KID ↓ InstantID + OMG 0.369 0.199 0.227 0.520 0.345 0.374 112.6 0.258 0.267 3.923 0.101 IPAdapter+ OMG 0.228 0.129 0.142 0.565 0.376 0.401 116.3 0.244 0.266 4.133 0.102 IPA-Face + OMG 0.310 0.166 0.188 0.515 0.341 0.367 115.4 0.249 0.267 4.021 0.102 Ours 0.421 0.298 0.317 0.630 0.548 0.560 60.65 0.506 0.273 4.237 0.093
Table 2: Top-1 preference rates from the user study on naturalness, face identity, body shape, and text-image correspondence in personalized multi-human scene generation.
Top 1 (%) Natural.↑ Face ID↑ Body shape↑ Text corr.↑ Textual Inversion 9.77 11.46 10.29 8.41 DreamBooth 15.4 11.13 13.33 13.01 InstantID + OMG 10.74 11.65 11.59 12.62 IPAdapter + OMG 12.75 11.78 12.62 14.3 IPA-Face + OMG 10.87 9.51 10.68 10.74 Ours 40.45 44.47 41.49 40.91
Occlusion Boundary Enhancer Network. Using this occlusion mask, we introduce an Occlusion Boundary Enhancer Network (OccNet) by training the model to generate images based solely on SMPLx depth edges within the occlusion mask. The SMPLx depth edges, defined as $\bm{e}_{\text{SMPLx}}=\left|\frac{\partial\bm{d}_{\text{SMPLx}}}{\partial x}\right|+\left|\frac{\partial\bm{d}_{\text{SMPLx}}}{\partial y}\right|>\tau,$ where $\tau$ is the edge threshold, highlight occlusion boundaries where depth changes abruptly. Although this information is limited due to the loss of 3D information inside the occluded regions, it serves as a crucial signal by highlighting the occlusion boundaries.
As shown in Fig. 3, we provide only the depth edges within the occluded region and train the model to generate images based on this sparse signal, which enables the model to focus more on generating the occluded region while preserving its boundaries, compared to training with the full conditioning. As illustrated in Fig. 4, we train OccNet $\mathcal{E}^{\text{Occ}}_{\omega}$ using the following objective: